# Alternating Direction Method of Multipliers AR AutoRally

### v1

@@
MPPI-Generic: A CUDA Library for Stochastic Optimization
@@ Abstract
This paper introduces a new C++/CUDA library for GPU-accelerated stochastic optimization called MPPI-Generic. It provides implementations of Model Predictive Path Integral control, Tube-Model Predictive Path Integral Control, and Robust Model Predictive Path Integral Control, and allows for these algorithms to be used across many pre-existing dynamics models and cost functions. Furthermore, researchers can create their own dynamics models or cost functions following our API definitions without needing to change the actual Model Predictive Path Integral Control code. Finally, we compare computational performance to other popular implementations of Model Predictive Path Integral Control over a variety of GPUs to show the real-time capabilities our library can allow for. Library code can be found at: https://acdslab.github.io/mppi-generic-website/
@@ I Introduction
As robotics and autonomy continue to grow, the choice of algorithms and methods used to plan and control of complex systems in these fields start to become more selective. These algorithms need to be able to handle the potentially complex dynamics of robotics system and allow representation for complex cost functions. They also need to be responsive to new data as environments and systems change, as in a self-driving scenario. Finally, the planning and control methods used need to be able to run in real-time so that the robot can continue to perform without unnecessary pauses.
@@
The approaches to control optimization can be delineated into two types of methods: gradient-based and sampling-based. Gradient-based methods such as iterative Linear Quadratic Regulator (iLQR) [1], Differential Dynamic Programming (DDP) [2] and Sequential Quadratic Programming (SQP) [3] generally have to rely on some restrictions to the dynamics and cost functions they are applied to such as being continuously differentiable. But in exchange for those restrictions, they can produce controls that minimize the cost function in a computationally-efficient manner. Sampling-based methods, such as Model Predictive Path Integral (MPPI) [4] or Cross-Entropy Method (CEM) [5], can relax these requirements and allow for arbitrary functions but come at the cost of requiring many samples to properly estimate the optimal control. Thus, we can see that while we gain more freedom in representation of functions, we pay for that in computational expense. One way to address this computational expense is to push the computation out of the CPU and onto a GPU. GPUs are designed to be highly parallelizable and for the majority of most sampling-based methods, we want to run the same computations on different inputs. By taking advantage of the GPU, we can make sampling-based methods usable in real time and also provide enough samples to get optimal solutions.
@@
MPPI utilizes the parallelization offered by GPUs. First introduced in [4], MPPI is a stochastic optimization Model Predictive Control (MPC) algorithm derived using the information theoretic dualities between relative entropy and free energy. Experiments in [4] included off-road navigation using the GT-Autorally vehicle [6] and demonstrated for the first time utilization of GPU for sampling-based MPC on real hardware. Follow-up works included the derivation of MPPI for the case of non-affine dynamics [7, 8] and extensions to multi-layer control architectures that incorporate the benefits of Differential Dynamic Programming to increase robustness to disturbances and model errors. These extensions include the Tube-based MPPI [9], Robust-MPPI (RMPPI) [10] architectures which consists of two layers of control.
@@
Since this first wave of publications on MPPI there has been a plethora of extensions and improvements including Tsallis-MPPI[11], Covariance Control MPPI [12], Constrained Covariance Steering MPPI [13], and Colored MPPI [14], Residual MPPI [15], Biased-MPPI [16], Multi-Modal MPPI [17], Risk-Aware MPPI [18], Spline-Interpolated Model Predictive MPPI [19], Stein Variational Guided MPPI [20], log-MPPI [21], and Smooth MPPI [22]. Besides the obvious use-case of MPC, alternative use-cases of MPPI can be found in the Model-Based Reinforcement Learning (MBRL) literature [23, 24, 25, 26, 27, 28]. In MBRL, MPPI is used to seek data from a simulation environment to learn the underlying value function and corresponding policy. Finally, while the majority of prior work on MPPI has direct applications to planning and control for systems in robotics and autonomy such as quad-rotors [29, 19], terrestrial [4, 14, 21, 30, 31], sea surface [32, 33] and underwater [34] vehicles, manipulators [35] and systems with multi-body dynamics [36], there are notable applications to other domains of science and engineering. These include control of HVAC systems [37, 38], chemical reactors [39], and pulse width modulation rectifiers [40].
@@
Given the far reaches of the MPPI algorithm, it is important to provide a solid and flexible computational framework from which researchers can make use of previous advancements to push their own work forwards. That framework needs to not only provide many built-in options to allow for testing on different platforms but also allow for researchers to develop new ideas on top of. Finally, that library needs to be able to support researchers all the way to hardware deployment with the ability to run in real time even on older hardware platforms.
@@
With this inspiration in mind, we introduce our controls and planning optimization library, MPPI-Generic. It is written in C++/CUDA and contains multiple dynamics and cost functions to allow for researchers to begin using them in complex robotics scenarios. In addition, it also allows for researchers to create their own dynamics or cost functions that can take advantage of the GPU-accelerated controllers. It provides implementations of MPPI[4], Tube-MPPI[9], and RMPPI [10] as each requires different GPU code to support. If desired, there is an API to allow for implementation of alternative sampling-based controllers. Initial versions of this library have already been used in a variety of hardware and software experiments [14, 32, 10, 41, 42]. To the best of our knowledge, this is the first MPPI library implementation to provide GPU-acceleration with real-time performance, multiple existing dynamics and cost functions, replaceable sampling distributions and controller algorithms, and extensibility options for researchers to create new components.
@@
  • We provide a NVIDIA GPU-accelerated Library for MPPI.
@@
The rest of the paper is organized as follows: In Section II, we provide the general problem formulation of a stochastic optimization problem and discuss the building blocks of the MPPI-Generic Library. In Section III, we discuss implementation details of our library and performance parameters available to the user to tweak. In Section IV, we show how to use the library through coding examples. In Section V, we show computational comparisons of our library against other implementations of MPPI. Finally, we state our conclusions in Section VI.
II Background
@@
${\bf x}_{t+1}$ $={\bf F}\left({\bf x}_{t},{\bf u}_{t}\right)$ (1)
${\bf y_{t}}$ $=G\left({\bf x}_{t},{\bf u}_{t}\right)$ (2)
where ${\bf x}\in\mathbb{R}^{n_{x}}$ is the state of dimension $n_{x}$, ${\bf u}\in\mathbb{R}^{n_{u}}$ is the control of dimension $n_{u}$, ${\bf y}\in\mathbb{R}^{n_{y}}$ is the observation of the system in dimension $n_{y}$, $T$ is the time horizon, $Y$ is an output trajectory $[{\bf y}_{1},{\bf y}_{2},...,{\bf y}_{T}]$, $U$ is a control trajectory $[{\bf u}_{0},{\bf u}_{1},...,{\bf u}_{T-1}]$, $\phi$ is the terminal cost, and ${\bf\ell}$ is the running cost.
Looking at the above equations, we can see that there is a minor difference from the typical nonlinear control setup, shown in Eq. 4, in that we have our cost function using the output, ${\bf y}_{t}$ instead of ${\bf x}_{t}$.
$\operatorname{\bf J}(X,U)$ $=\phi({\bf x}_{T})+\sum_{t=0}^{T-1}{\bf\ell}\left({\bf x}_{t},{\bf u}_{t}\right)$ (4)
For the vast majority of systems, ${\bf y}_{t}$ can just be the true state, i.e. ${\bf y}_{t}={\bf x}_{t}$ but we have seen that separating the two can be important for computational reasons. It allows reuse of computationally-heavy calculations required for dynamics updates such as wheel position and forces in the cost function. Having this split between ${\bf x}_{t}$ and ${\bf y}_{t}$ is one that allows us to improve the efficiency of the real-time code while not having to do unnecessary computations such as a Jacobian with respect to every wheel position.
It is important to note that the MPPI-Generic dynamics assumes Euler integration by default but can be modified if desired by developing a custom dynamics class.
@@ II-A The MPPI Algorithm
MPPI is a stochastic optimal control algorithm that minimizes Eq. 4 through the use of sampling. We start by sampling control trajectories, running each trajectory through the dynamics in Eq. 1 to create a corresponding state trajectory, and then evaluating each state and control trajectory through the cost function. Each trajectory’s cost is then run through the exponential transform,
$S({\bf J};\lambda)=\exp\left(-\frac{1}{\lambda}{\bf J}\right),$ (5)
@@
where $\lambda$ is the inverse temperature. Finally, a weighted average of the trajectories is conducted to produce the optimal control trajectory. The update law for $\mathcal{U}^{*}(t)$, the optimal trajectory at time $t$, ends up looking like
$\mathcal{U}^{*}_{t}$ $=\sum_{m=1}^{M}\frac{\exp\left(-\frac{1}{\lambda}{\bf J}\left(X^{m},V^{m}\right)\right){\bf v}^{m}_{t}}{\sum_{j=1}^{M}\exp\left(-\frac{1}{\lambda}{\bf J}\left(X^{j},V^{j}\right)\right)}$ (6)
$={\bf u}_{t}+\sum_{m=1}^{M}\frac{\exp\left(-\frac{1}{\lambda}{\bf J}\left(X^{m},V^{m}\right)\right)\epsilon^{m}_{t}}{\sum_{j=1}^{M}\exp\left(-\frac{1}{\lambda}{\bf J}\left(X^{j},V^{j}\right)\right)},$ (7)
@@
where $V^{m}$ is the $m$-th sampled control trajectory, ${\bf v}^{m}_{t}={\bf u}_{t}+\epsilon^{m}_{t}$ is the control from the $m$-th sampled trajectory at time $t$ sampled around the previous optimal control, ${\bf u}_{t}$, with $\epsilon^{m}_{t}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. Sampling in the control space ensures that the trajectories are dynamically feasible and allows us to use non-differentiable dynamics and cost functions. Pseudo code for the algorithm is shown in Algorithm 1.
@@
Given :  ${\bf F}\left(\cdot,\cdot\right)$, $G\left(\cdot,\cdot\right)$ $\ell\left(\cdot,\cdot\right)$, $\phi\left(\cdot\right)$ $M$, $I$, $T$, $\lambda$, $\sigma$: System dynamics, system observation, running cost, terminal cost, num. samples, num. iterations, time horizon, temperature, standard deviations;

### v2

@@
MPPI-Generic: A CUDA Library for Stochastic Trajectory Optimization
@@ Abstract
This paper introduces a new C++/CUDA library for GPU-accelerated stochastic trajectory optimization called MPPI-Generic. It provides implementations of Model Predictive Path Integral control, Tube-Model Predictive Path Integral Control, and Robust Model Predictive Path Integral Control, and allows for these algorithms to be used across many pre-existing dynamics models and cost functions. Furthermore, researchers can create their own dynamics models or cost functions following our API definitions without needing to change the actual Model Predictive Path Integral Control code. Finally, we compare computational performance to other popular implementations of Model Predictive Path Integral Control over a variety of GPUs to show the real-time capabilities our library can allow for. Our library can be found at: https://acdslab.github.io/mppi-generic-website/
@@ I Introduction
As robotics and autonomy continue to grow, the choice of algorithms and methods used to plan and control complex systems in these fields start to become more selective. These algorithms need to be able to handle the potentially intricate dynamics found in robotic systems and allow for complex cost function representations. They also must be responsive to new data as environments and systems change. Finally, the planning and control methods used should run in real-time so that the robot can continue to perform without unnecessary pauses. We can see these requirements show themselves in a self-driving car example. The dynamics need to be aware of changing road conditions while the cost function has to capture multiple goals such as avoiding other cars, following traffic laws, and getting to the destination. The planner has to be fast enough to react to other drivers stopping suddenly or debris on the road.
@@
The approaches to Model Predictive Control (MPC) optimization can be delineated into two types of methods: gradient-based and sampling-based. Gradient-based methods such as iterative Linear Quadratic Regulator (iLQR) [1], Differential Dynamic Programming (DDP) [2] and Sequential Quadratic Programming (SQP) [3] generally rely on restrictions to the dynamics and cost functions such as being continuously differentiable. But in exchange for those restrictions, they can produce controls that minimize the cost function in a computationally-efficient manner. Sampling-based methods, such as Model Predictive Path Integral (MPPI) [4] or Cross-Entropy Method (CEM) [5], can relax these requirements and allow for arbitrary functions but come at the cost of requiring many samples to properly estimate the optimal control. These arbitrary functions remove the need for convexification or smoothing which can cause unnecessary conservatism. One way to address this computational expense is to push the computation out of the CPU and onto a GPU, where the parallelization of sampling can be better utilized. By taking advantage of the GPU, we make sampling-based methods usable in real-time and also provide enough samples to get optimal solutions.
@@
First introduced in [4], MPPI is a stochastic MPC algorithm derived using the information theoretic dualities between relative entropy and free energy. Experiments in [4] included off-road navigation using the GT-Autorally vehicle [6] and demonstrated for the first time utilization of GPU for sampling-based MPC on real hardware. Follow-up works included the derivation of MPPI for the case of non-affine dynamics [7, 8] and extensions to multi-layer control architectures that incorporate the benefits of iLQR to increase robustness to disturbances and model errors. These extensions include the Tube-based MPPI [9], Robust-MPPI (RMPPI) [10, 11] architectures which consists of two layers of control. Further discussion of extensions to MPPI is left to Section II-B.
@@
Besides the obvious use-case of MPC, alternative use-cases of MPPI can be found in the Model-Based Reinforcement Learning (MBRL) literature [12, 13, 14, 15, 16, 17]. In MBRL, MPPI is used to seek data from a simulation environment to learn the underlying value function and corresponding policy. Finally, while the majority of prior work on MPPI has direct applications to planning and control for systems in robotics and autonomy such as quad-rotors [18, 19], terrestrial [4, 20, 21, 22, 23], sea surface [24, 25] and underwater [26] vehicles, manipulators [27], and systems with multi-body dynamics [28], there are notable applications to other domains of science and engineering. These include control of HVAC systems [29, 30], chemical reactors [31], and pulse width modulation rectifiers [32].
@@
Given the far reaches of the MPPI algorithm, it is important to provide a solid and flexible computational framework from which researchers can make use of previous advancements to push their own work forwards. That framework needs to not only provide many built-in options to allow for testing on different platforms but also allow for researchers to develop new ideas on top of. Finally, that library needs to be able to support researchers all the way to hardware deployment with the ability to run in real-time even on older hardware platforms.
@@
With this inspiration in mind, we introduce our controls and planning optimization library, MPPI-Generic. It is written in C++/CUDA and contains multiple dynamics and cost functions to allow for researchers to begin using them in complex robotics scenarios. In addition, it also allows for researchers to create their own dynamics or cost functions that take advantage of the GPU-accelerated controllers. It provides implementations of MPPI[4], Tube-MPPI[9], and RMPPI [11] as well as an API for implementations of new sampling-based controller. Initial versions of this library have already been used in a variety of hardware and software experiments [20, 24, 11, 33, 34]. To the best of our knowledge, this is the first MPPI library implementation to provide GPU-acceleration with real-time performance, multiple existing dynamics and cost functions, replaceable sampling distributions and controller algorithms, and extensibility options for researchers to create new components.
@@
  • We provide a NVIDIA GPU-accelerated Library [^1] Library code is located at https://github.com/ACDSLab/MPPI-Generic for MPPI.
@@
The rest of the paper is organized as follows: In Section II, we provide the general problem formulation of a stochastic trajectory optimization problem, introduce various extensions to MPPI, and discuss the building blocks of the MPPI-Generic Library. In Section III, we discuss implementation details of our library and performance parameters available to the user to tweak. In Section IV, we show how to use the library through coding examples. In Section V, we show computational comparisons of our library against other implementations of MPPI. Finally, we state our conclusions in Section VI.
II Background on Stochastic Trajectory Optimization
@@
${\bf x}_{t+1}$ $={\bf F}\left({\bf x}_{t},{\bf u}_{t}+{\bf n}_{t}\right)$ (1)
$\operatorname{\bf J}(X,U)$ $=\phi({\bf x}_{T})+\sum_{t=0}^{T-1}{\bf\ell}\left({\bf x}_{t},{\bf u}_{t}\right),$ (2)
where ${\bf x}\in\mathbb{R}^{n_{x}}$ is the state, ${\bf u}\in\mathbb{R}^{n_{u}}$ is the control, ${\bf n}$ is assumed to be zero-mean Gaussian disturbances with variance $\Sigma$ in the control channel, $T$ is the time horizon, $X$ is a state trajectory $[{\bf x}_{0},{\bf x}_{1},...,{\bf x}_{T}]$, $U$ is a control trajectory $[{\bf u}_{0},{\bf u}_{1},...,{\bf u}_{T-1}]$, $\phi$ is the terminal cost, and ${\bf\ell}$ is the running cost. It is important to note that existing MPPI-Generic dynamics assume Euler integration by default but can be modified if desired by developing a custom dynamics class.
@@ II-A The MPPI Algorithm
While MPPI was originally derived from a path-integral approach, there have been other derivations that lead to slightly different update rules. We will briefly go over the information-theoretic derivation [8] to underpin the specific update rule we use in MPPI-Generic. We also briefly discuss other variations of MPPI that are included in this library.
II-A1 Information Theoretic Derivation
We shall follow the information-theoretic derivation from [8] with some changes. In [8], the running cost definition was only based on the state trajectory and the authors showed how a control cost would emerge naturally from the importance sampling scheme. However, the derivation does not require this limited running cost definition, so we generalize our running cost to $\ell({\bf x}_{t},{\bf u}_{t})$. We use the following shorthand to denote the cost for a control trajectory $U$ applied to the starting state, ${\bf x}_{0}$,
$\begin{split}\mathcal{J}\left(U\right)&:={\bf J}(X,U)\\ \text{s.t. }&{\bf x}_{t+1}={\bf F}\left({\bf x}_{t},{\bf u}_{t}\right)\end{split}$ (3)
Our goal is to find a control trajectory, $U^{*}$, that minimizes Eq. 2 through the use of sampling. We start by defining the free energy of our system as
$\mathcal{F}\left(\mathcal{J},\mathbb{P},\lambda;{\bf x}_{0}\right)$ $=-\lambda\ln\left(\mathbb{E}_{V\sim\mathbb{P}}\left[\exp\left(-\frac{1}{\lambda}\mathcal{J}\left(V\right)\right)\right]\right)$ (4)
where ${\bf x}_{0}$ is the initial state, $\lambda\in\mathbb{R}^{+}$ is the inverse temperature, and $V$ is the control trajectory sampled from some base distribution $\mathbb{P}$ with Probability Density Function (PDF) $p$. The base distribution $\mathbb{P}$ can be any distribution, making it potentially computationally intractable. Introducing a new distribution $\mathbb{Q}$ that is absolutely continuous with $\mathbb{P}$, we use importance sampling and Jensen’s Inequality to get the following,
$\mathcal{F}\left(\mathcal{J},\mathbb{P},\lambda;{\bf x}_{0}\right)$ $\leq\mathbb{E}_{V\sim\mathbb{Q}}\left[\mathcal{J}\left(V\right)\right]+\lambda\mathbb{KL}\left(\mathbb{Q}\ ||\ \mathbb{P}\right)$ (5)
where $\mathbb{KL}\left(\cdot\ ||\ \cdot\right)$ is the Kullback Leibler (KL) divergence between $\mathbb{Q}$ and $\mathbb{P}$. This introduces an upper bound on our free energy when we sample from $\mathbb{Q}$ instead of the original $\mathbb{P}$; this upper bound becomes a strict equality if $\mathbb{Q}^{*}$ has the following density:
$q^{*}(V)$ $=\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\mathcal{J}\left(V\right)\right)p(V)$ (6)
$\eta$ $=\mathbb{E}_{\hat{V}\sim\mathbb{P}}\left[\exp\left(-\frac{1}{\lambda}\mathcal{J}(\hat{V})\right)\right]$ (7)
While $\mathbb{Q}^{*}$ is an optimal distribution that minimizes the free energy bound, the optimal PDF, $q^{*}(V)$, still relies on the PDF of the original distribution $\mathbb{P}$, meaning it still might be computationally intractable. Thus, we introduce a third distribution $\mathbb{S}_{\theta}$ that has controllable parameters $\theta$ and minimizes the KL divergence between $\mathbb{S}_{\theta}$ and $\mathbb{Q}^{*}$,
$\theta^{*}$ $=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{KL}\left(\mathbb{Q}^{*}\ ||\ \mathbb{S}_{\theta}\right)$ (8)
$=\operatorname*{arg\,max}_{\theta\in\Theta}\mathbb{E}_{V\sim\mathbb{Q}^{*}}\left[\ln\left(s\left(V|\theta\right)\right)\right].$ (9)
Eq. 9 gives us an update rule for any probability distribution with parameters to match the optimal distribution. Leaving this aside for the moment, we show how to calculate the optimal control sequence by sampling from $\mathbb{S}_{\theta}$. From there, we will show how the update rule and the optimal control sequence can relate to each other depending on the choice of distributional family, $\mathbb{S}_{\theta}$. For the optimal control to use at time $t$, we shall look to the mean of the optimal distribution $\mathbb{Q}^{*}$,
${\bf u}_{t}^{*}$ $=\mathbb{E}_{V\sim\mathbb{Q}^{*}}\left[{\bf v}_{t}\right]$ (10)
[…]
Using importance sampling to sample from $\mathbb{S}_{\theta}$ as well as Eq. 6 gives the following,
[…]
$=\mathbb{E}_{V\sim\mathbb{S}_{\theta}}\left[w(V){\bf v}_{t}\right]$ (13)
$w(V)$ $=\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\mathcal{J}\left(V\right)\right)\frac{p\left(V\right)}{s\left(V|\theta\right)}$ (14)
The only term in $w(V)$ that is unaccounted for is the importance sampling weight of $\mathbb{P}$ and $\mathbb{S}_{\theta}$.
In [8], the base and parameterized distributions were both assumed to be Gaussian distributions with constant variance across time, $\Sigma$. The mean trajectory of $\mathbb{P}$ was assumed to be ${\bf 0}_{n_{u}\times T}$ while the mean trajectory of $\mathbb{S}_{\theta}$ was the parameter to optimize, i.e. $\theta={\bf U}^{*}=\left\{{\bf u}^{*}_{t}\right\}_{t=0}^{T-1}$. This led to the specific form of
$w(V)$ $=\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\mathcal{J}(V)-\sum_{t=0}^{T-1}\big{(}{\bf v}_{t}-\frac{1}{2}{\bf u}^{*}_{t}\big{)}^{\top}\Sigma^{-1}{\bf u}^{*}_{t}\right)$ (15)
and also meant that the solutions to Eqs. 9 and 13 were the same.
II-B Other MPPI Modifications
Beyond the derivations of MPPI, there have been other modifications that address various limiations of MPPI. A Tube-based MPPI controller [9] was created in order to improve robustness to state disturbances. It made use of a tracking controller to track the real system back to a nominal system that ignored state disturbances resulting in large costs. The nominal system would use an initial state produced from the dynamics equation in Eq. 1 for a new iteration of MPPI instead of the real system’s state estimate. Both the real and the nominal trajectories are calculated using MPPI while the tracking controller was iLQR. In this setup, the tracking controller would always be the one sending controls to the system and as MPPI was not aware of the tracking controller, it could end up fighting against the tracking controller. In order to address that, RMPPI was developed in [10, 11], which applied the tracking controller feedback within the samples MPPI used. RMPPI also chose the initial state for the nominal system using a constrained optimization problem that tried to keep the nominal state as close to the real state without causing the resulting trajectory to have a cost larger than a given cost threshold $\alpha$. Choosing the nominal state in this way ensured there was an upper bound on how quickly the optimal trajectory’s cost could grow due to disturbance. However, this improved robustness to state disturbances can struggle when the cost function itself changes over time due to new information. Our library contains implementations of these algorithmic improvements to MPPI as different controllers are the best choice in different scenarios.
Since this first wave of publications on MPPI, there has been a plethora of extensions and improvements. These include sampling efficiency improvements such as Covariance Control MPPI [35], Colored MPPI [20], Residual MPPI [36], Biased-MPPI [37], Spline-Interpolated Model Predictive MPPI [19], Stein Variational Guided MPPI [38], log-MPPI [21], CoVO-MPC [39], U-MPPI [40], o-MPPI [41], and Smooth MPPI [42]; there have also been robustness improvements such as Tsallis-MPPI[43], Constrained Covariance Steering MPPI [44], Multi-Modal MPPI [45], and Risk-Aware MPPI [46]. Some of the sampling efficiency methods have also been implemented, shown in Section IV-C3, but the rapid development of the community makes it nearly impossible to keep up.
II-C MPPI-Generic’s MPPI Implementation
In practice, MPPI-Generic uses the update rule, Eq. 13, with Monte-Carlo sampling to find the optimal control sequence. The algorithm starts by sampling control trajectories, running each trajectory through the dynamics in Eq. 1 to create a corresponding state trajectory, and evaluating each state and control trajectory through the cost function. We provide the option to calculate an importance sampling ratio calculated for the specific sampling distribution like in Eq. 14 or to disable it,
$\mathcal{I}\left(V\right)$ $=\beta\ln\left(\frac{p\left(V\right)}{s\left(V|\theta\right)}\right)$ (16)
@@ II-C MPPI-Generic’s MPPI Implementation
where $\beta\in\left\{0,1\right\}$. The option to disable the importance sampling weight comes from experimentation, such as on the AutoRally platform [6], where we have seen better control trajectories without it. Finally, a weighted average of the trajectories is conducted to produce the optimal control trajectory. The update law for $\mathcal{U}^{*}_{t}$, the optimal control at time $t$, ends up looking like
$\mathcal{U}^{*}_{t}$ $=\sum_{m=1}^{M}\frac{\exp\left(-\frac{1}{\lambda}\left(\mathcal{J}\left(V^{m}\right)-\lambda\mathcal{I}\left(V^{m}\right)-\rho\right)\right){\bf v}^{m}_{t}}{\sum_{j=1}^{M}\exp\left(-\frac{1}{\lambda}\left(\mathcal{J}\left(V^{m}\right)-\lambda\mathcal{I}\left(V^{m}\right)-\rho\right)\right)}$ (17)
$\rho$ $=\min_{m\in\left[1,M\right]}\left(\mathcal{J}\left(V^{m}\right)-\lambda\mathcal{I}\left(V^{m}\right)\right)$ (18)
@@
where $V^{m}$ is the $m$-th sampled control trajectory, ${\bf v}^{m}_{t}$ is the control from the $m$-th sampled trajectory at time $t$, and $\rho$ is the minimum sampled cost. When using a Gaussian distribution, the sampled control ${\bf v}^{m}_{t}={\bf u}_{t}+\epsilon^{m}_{t}$ is centered around the previous optimal control ${\bf u}_{t}$ and has noise $\epsilon^{m}_{t}\sim\mathcal{N}\left(0,\Sigma\right)$; other distributions would have different ways to draw samples and alternative $\mathcal{I}\left(V\right)$ equations. We use $\rho$ as in [8] to shift the range of the exponentiated costs to limit numerical stability issues. Sampling in the control space ensures that the trajectories are dynamically feasible and allows us to use non-differentiable dynamics and cost functions. Pseudo-code for the algorithm is shown in Algorithm 1.
@@
Given :  ${\bf F}\left(\cdot,\cdot\right)$, $G\left(\cdot,\cdot\right)$ $\ell\left(\cdot,\cdot\right)$, $\phi\left(\cdot\right)$, $\mathcal{I}\left(\cdot\right)$, $M$, $I$, $T$, $\lambda$, $\Sigma$: System dynamics, system observation, running cost, terminal cost, importance sampling ratio, num. samples, num. iterations, time horizon, temperature, covariance;